OpenStudy (anonymous):

Will give Medal/Fan :) A=(pi)r^2 Find the inverse of function(9)

OpenStudy (anonymous):

@jessiegonzales @vzfreakz could either of you possible help? :(

OpenStudy (anonymous):

3/pi^2 is what i got

OpenStudy (anonymous):

Quite forgot how to solve this. Function 9, hmm. Where would the 9 be, in the equation?

OpenStudy (anonymous):

f−1(9)

OpenStudy (anonymous):

f-1(9)

OpenStudy (anonymous):

the equation equals 9

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

Give me your equation in one. Just to be sure you're correct.

OpenStudy (anonymous):

in one?

OpenStudy (anonymous):

i put it up top

OpenStudy (anonymous):

Hm.. it's been a while my friend. Can't get it in my head about, sorry, try to ask someone else.

OpenStudy (anonymous):

ok thanks anyways

OpenStudy (anonymous):

any suggestions

OpenStudy (anonymous):

Yes, but hopefully they comment back :)

OpenStudy (anonymous):

@ganeshie8 @IMStuck bros, I really need your help. A fellow mate needs help with his assignment. Please help him real quick. Thanks!

OpenStudy (anonymous):

thanks buddy

OpenStudy (anonymous):

You can post another one if you like.

OpenStudy (anonymous):

ill tag you if i need more help i have 5 more, dont want to keep you waiting. i gave you the medal/fan though, thanks

OpenStudy (anonymous):

Thanks, really appreciate it. But I don't think I did enough to help you out.

OpenStudy (anonymous):

still there?

ganeshie8 (ganeshie8):

A=(pi)r^2 I think "r" needs to be isolated here

ganeshie8 (ganeshie8):

\[\large r = \sqrt{ \dfrac{A}{\pi}}\]